# On an Optimal Stopping Problem with a Discontinuous Reward

### v2

@@ 4.3. Surrender and Continuation Premium Representation
In the following, we derive two representations for the value function. The first one is akin to the early exercise premium representation (or integral representation) in the American option terminology.
To the authors’ knowledge, the second representation is new to the literature on variable annuities and American options. The particular form of the reward function involved in variable annuity pricing allows the decomposition of the value function in two terms: the current value of the reward process and an integral term which increases only when the sub-account process is in the continuation region. Therefore, we call this second term the continuation premium.
@@
Let Assumptions 4.1, 4.2 and 4.3 hold. Then, the value function can be written as
@@
where $h:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is the present value of the maturity benefit, as defined in (5), and $e:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ denotes the early surrender premium
@@
$\int_{t}^{T}e^{-r(s-t)}\sigma F_{s}^{t,x}v_{x}(s,\,F_{s}^{t,x})\,\mathrm{d}W_{s}$ is a martingale since $|v_{x}(t,x)|\leq 1$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, as per Proposition 4.1 and Corollary 4.2, and ${\mathbb{E}[\int_{0}^{T}F_{s}^{2}\,\mathrm{d}s]<\infty}$. Thus, taking the expectation on both sides of (29) yields
@@
The proof of Proposition 4.4 requires the value function to be smooth enough to apply Itô’s formula for generalized derivatives. However, when the reward function is continuous, a similar characterization of the value function can be obtained in a very general context, see Jacka (1993), Theorem 6 and the remark on p.337. See also Terenzi (2019), Proposition 2.4.10, for an application of Jacka (1993)’s results to American put option pricing under the Heston model. The results of Jacka (1993) can be used with our continuous reward representation (14). However, as observed in Example 3.3 and Remark 3.5, the original problem in (7) and its modified version in (14) result in the same value function (Theorem 3.4), but can lead to different surrender regions and optimal surrender boundary. Therefore, to derive the results of Proposition 4.4 in a more general market model using Jacka (1993) and Theorem 3.4 of this paper, one first needs to show the equivalence of the two optimal stopping problems, (7) and (14), in this general setting. This is left as future research.
Remark 4.10.
If $L(t)<0$ for all $t\in[0,T)$, the surrender region has the particular shape $\mathcal{S}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|x\geq b(t)\}$ for some $b(t)\geq Ge^{-r(T-t)}$. This will be shown in the next section in Proposition 4.8. Under this assumption, the early surrender premium becomes
@@
The next results provide a representation of the value function in terms of the loss incurred for not exercising optimally.
@@
The value function can be written as
@@
where ${f:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}}$ denotes the continuation premium defined as
@@
The result presented in Proposition 4.5 is very interesting since it decomposes the value function in terms of the immediate surrender value plus a term representing the value of holding on to the contract. This term, which we coin the continuation premium, is the sum of the financial guarantee at maturity (or the put option) when the contract is held until $T$ and an integral term equal to the value added by keeping the contract until it is optimal to surrender. Theorem 4.3 below shows that with some additional assumptions on the fee and the surrender functions, the dependence of $f$ on $\tau_{t}^{x}$ can be removed. The representation of the continuation premium in (40) can be particularly helpful to develop numerical methods for approximating the value of a variable annuity contract, since it uses additional information on the shape of the continuation region. The results of Theorem 4.3 rely on the following Lemma.
@@
Suppose Assumptions 4.1, 4.2 and 4.3 hold. For all $(t,x)\in[0,T]\times\mathbb{R}_{+}$,
@@
Note that additional assumptions are required in Lemma 4.3 since the proof makes use of the surrender premium representation (27) in Proposition 4.4.
Theorem 4.3.
Suppose Assumptions 4.1, 4.2 and 4.3 hold. The continuation premium $f(t,x)$ of Proposition 4.5 can be written as
Proof.
@@
If $L(t)<0$ for all $t\in[0,T)$, the continuation region is given by ${\mathcal{C}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|x<b(t)\}}$ for some $b(t)\geq Ge^{-r(T-t)}$, see Proposition 4.8. Under this assumption, the continuation premium in (40) becomes
4.4. Characterization of the Surrender Region
The goal of this section is to study the properties of the surrender and continuation regions. Under certain conditions on the fee and surrender charge functions, we can fully characterize the two regions.
The results below are inspired by the work of Villeneuve (1999), which we adapted to the time-dependent and discontinuous reward function considered in (7). As pointed out by Villeneuve (1999), Remark 2.1, adapting some of their results to a time-dependent payoff is not trivial. The next two results characterize the non-emptiness of the surrender region. We first recall that, under Assumption 4.2 , the function $L$, defined in (22), becomes $L(t)=g_{t}(t)-c(t)g(t).$
@@ 4.4. Characterization of the Surrender Region
    For each $t\in[0,T)$, if $\mathcal{S}_{t}\neq\emptyset$, then $L(t)\leq 0.$
@@
    Further let Assumptions 4.1 and 4.3 hold. If $L(t)<0$ for all $t\in[0,T)$, then $\mathcal{S}\neq\emptyset$ and $\mathcal{C}\neq\emptyset$.
@@
    Observe that for $x\in\mathcal{S}_{t}$, $v(t,x)=xg(t)$, so that $xL(t)=\mathcal{L}_{t}v+v_{t}-rv$. Now recall from Proposition 4.2 that under Assumption 4.2, $\mathcal{L}_{t}v+v_{t}-rv\leq 0$ for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$. The result follows since $x>0$ in $\mathcal{S}_{t}$.
@@
    We show that if $L(t)<0$ for all $t\in[0,T)$ then $\mathcal{S}\neq\emptyset$ and $\mathcal{C}\neq\emptyset$. Using the continuation premium representation (40), we have
@@
    for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. We want to show that there exists some $(t_{1},x_{1})\in[0,T)\times\mathbb{R}_{+}$ satisfying $v(t_{1},x_{1})=\varphi(t_{1},x_{1})$ and some $(t_{2},x_{2})\in[0,T)\times\mathbb{R}_{+}$ satisfying ${v(t_{2},x_{2})>\varphi(t_{2},x_{2})}$. We proceed by contradiction. First, suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{S}=\emptyset$. Fix $t\in[0,T)$. Since $\mathcal{S}=\emptyset$, we have $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, so that (42) becomes
@@
    Observe now that $L(s)=g_{t}(s)-c(s)g(s)<0\Leftrightarrow\frac{\,\mathrm{d}\ln g(s)}{\,\mathrm{d}s}<c(s)$ for all $s\in[0,T)$. By integrating from $t$ to $T$ on both sides of the inequality and using the fact that $g(T)=1$, we find that
@@
We now state a local version of Proposition 4.6 (ii).
Lemma 4.4.
Let Assumptions 4.1, 4.2 and 4.3 hold. If $L(t)<0$ for all $t\in[0,T)$, then $\cup_{{t_{1}\leq t<t_{2}}}\mathcal{S}_{t}\neq\emptyset$ for any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$.
@@
We proceed by contradiction. Fix $0\leq t_{1}<t_{2}<T$. Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\cup_{t_{1}\leq t<t_{2}}S_{t}=\emptyset$ This implies $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, so that $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Now fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Using the continuation premium representation (40), we have that
$\begin{split}v(t,x)-\varphi(t,x)&=\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\\ &\quad+\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\right]\,\mathrm{d}s>0,\end{split}$
so that,
$\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]$ $>-\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\right]\,\mathrm{d}s$
$\geq-\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\right]\,\mathrm{d}s$
(46) $=-x\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s$
Note that the function on the right-hand side is positive, equal to 0 when $x=0$, strictly increasing in $x$ (since $L(t)<0$ by assumption) and approaches infinity as $x\rightarrow\infty$. Then, using arguments similar to those of the proof of Proposition 4.6 (ii), there must exist ${y\in\mathbb{R}_{+}}$ such that for all ${x\geq y}$, ${\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\leq-x\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s}$, contradicting the strict inequality in (46) which must hold for all ${x\in\mathbb{R}_{+}}$. We conclude that $\mathcal{S}_{t}\neq\emptyset$.
The next results characterize the emptiness of the surrender region.
Proposition 4.7.
Let Assumption 4.2 hold.
@@
    For each $t\in[0,T)$, if $L(t)>0$, then $\mathcal{S}_{t}=\emptyset$.
@@
    Further let Assumptions 4.1 and 4.3 hold. If $\mathcal{S}=\emptyset$, then $g(t)-e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}\leq 0$ for all $t\in[0,T)$.
Proof.
The first statement is simply the contrapositive of Proposition 4.6 (i). For the second statement, since $\mathcal{S}=\emptyset$, $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Then, for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ implies
(48) $g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0$
for all $t\in[0,T)$. ∎
Remark 4.12.
Proposition 4.7 (ii) states that an empty surrender region results in ${g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0}$ being satisfied for all $t\in[0,T)$. This result is weaker than finding $L(t)\geq 0$ for all $t\in[0,T)$. Therefore, we only have a partial converse statement linking the two parts of Proposition 4.7. Indeed, if $L(t)\geq 0$ for all $t\in[0,T)$, then $g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0$ for all $t\in[0,T)$. To see this, fix $t\in[0,T)$, and observe that $L(t)\geq 0$ implies that $\frac{\,\mathrm{d}\ln g(t)}{\,\mathrm{d}t}\geq c(t)$, from which we get
$g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0.$
However, the inverse statement does not hold; that is, $g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0$ for all $t\in[0,T)$ does not necessarily imply $L(t)\geq 0$ for all $t\in[0,T)$, since the integral of $L(t)$ from 00 to $T$ can be positive even when $L(s)$ is negative for some $s\in[0,T]$.
Remark 4.13.
The first statements of Propositions 4.6 and 4.7 can apply to surrender charge functions depending on both time and the value of the underlying process. Indeed, the proof relies on the results of Proposition 4.2 which only requires the fee function, and not the surrender charge, to depend exclusively on time, see Remark 4.7 for details.
Proposition 4.6 (ii) and Proposition 4.7 (i), state that the surrender region being empty or not only depends on the sign of $L(t)$. We believe that the results of Proposition 4.6 (ii) can be made more local. That is, we conjecture that as soon as $L(t)<0$ for some $t\in[0,T)$, the $t$-section of the surrender region is not empty. This is what we observe in numerical examples, see Section 5. Hence, when $L(t)$ changes sign in $[0,T)$, the boundary of the continuation region, $\partial\mathcal{C}$, is discontinuous. This property of the optimal surrender boundary sets the optimal stopping problem involved in variable annuities pricing apart from the pricing of standard American call and put options.
The next result relaxes Assumption 4.2 in Proposition 4.7 (i) and gives a local version of the statement on the emptiness of the surrender region.
Lemma 4.5.
For any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, then $\cup_{{t_{1}\leq t<t_{2}}}\mathcal{S}_{t}=\emptyset$.
Proof.
First, we observe that $\cup_{{t_{1}\leq t<t_{2}}}\mathcal{S}_{t}=\emptyset$ implies ${\cup_{{t_{1}\leq t<t_{2}}}\mathcal{C}_{t}=[t_{1},t_{2})\times\mathbb{R}_{+}}$. Thus, we need show that $v(t,x)>\varphi(t,x)$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$ and recall that from the proof of Proposition 4.5 that for any $0<t<s\leq T$
@@
for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, which completes the proof. ∎

### v3

@@ 4.3. Surrender and Continuation Premium Representation
In this section and throughout the rest of the paper, derivatives should be understood as distributional.
Below, we derive two representations for the value function. The first one is akin to the early exercise premium representation (or integral representation) in the American option terminology.
To the authors’ knowledge, the second representation is new to the literature on variable annuities and American options. The particular form of the reward function involved in variable annuity pricing allows the decomposition of the value function in two terms: the current value of the reward process and an integral term which increases only when the account process is in the continuation region. Therefore, we call this second term the continuation premium.
@@
The value function (7) can be written as
@@
where $h:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is the present value of the maturity benefit, as defined in (5), and $e:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}$ denotes the early surrender premium
@@
$\int_{t}^{T}e^{-r(s-t)}\sigma F_{s}^{t,x}v_{x}(s,\,F_{s}^{t,x})\,\mathrm{d}W_{s}$ is a martingale since $|v_{x}(t,x)|\leq 1$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, as per Proposition 4.1 and Corollary 4.2, and ${\mathbb{E}[\int_{0}^{T}F_{s}^{2}\,\mathrm{d}s]<\infty}$. Thus, taking the expectation on both sides of (LABEL:eqSurrPrem1) yields
@@
If $L(t)<0$ for all $t\in[0,T)$, the surrender region has the particular shape $\mathcal{S}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|x\geq b(t)\}$ for some $b(t)\geq Ge^{-r(T-t)}$. This will be shown in the next section in Theorem 4.3. Under this assumption, the early surrender premium becomes
@@
The value function (7) can be written as
@@
where ${f:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}}$ is the continuation premium given by
The proof of Proposition 4.5 is presented at the end of the section; it relies on the results below.
Remark 4.13.
If $L(t)<0$ for all $t\in[0,T)$, the continuation region is given by ${\mathcal{C}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|x<b(t)\}}$ for some $b(t)\geq Ge^{-r(T-t)}$, see Theorem 4.3. Under this assumption, the continuation premium in (28) becomes
Proposition 4.5 decomposes the value function in terms of the immediate surrender value plus a term representing the value of holding on to the contract. This term, which we coin the continuation premium, is the sum of the financial guarantee at maturity (or the put option) when the contract is held until $T$ and an integral term equal to the value added by keeping the contract until it is optimal to surrender. The representation of the continuation premium in (28) can be particularly helpful to develop numerical methods for approximating the value of a variable annuity contract, since it uses additional information on the shape of the continuation region.
It is also possible to decompose the value function into the surrender benefit and the continuation premium in a more general setting, where Assumptions 4.1 and 4.2 do not need to hold. However, in this case, the continuation premium is written in terms of $\tau_{t}^{x}$, which can make it more complicated to calculate. This result is presented in Lemma 4.3 below and will be used to prove Proposition 4.5.
Lemma 4.3.
The value function (7) can be written as
(30) $v(t,x)=xg(t,x)+f(t,x),$
where ${f:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}}$ is the continuation premium given by
@@
Remark 4.14.
As stated above, and as can be seen from its proof, Lemma 4.3 holds even when Assumptions 4.1 and 4.2 are not satisfied.
The proof of Proposition 4.5 also relies on the following Lemma, which is used to remove the dependence of the continuation premium on $\tau_{t}^{x}$.
@@
For all $(t,x)\in[0,T]\times\mathbb{R}_{+}$,
@@
We note that Lemma 4.4 requires Assumptions 4.1 and 4.2 to hold since the proof makes use of the surrender premium representation (23) in Proposition 4.4.
Lemmas 4.3 and 4.4 can now be used to prove Proposition 4.5.
Proof of Proposition 4.5.
@@
4.4. Characterization of the Surrender Region
In this section, we study the impact of the fee and surrender charge functions on the shape of the surrender (and continuation) region. The main results of this section characterize specific types of surrender regions and are summarized in Theorem 4.3 below. The rest of the section contains auxiliary results that are used in the proof of Theorem 4.3 and that are of interest on their own. The proof of Theorem 4.3 is presented at the end of the section.
Henceforth, the $t$-section of the surrender (resp. continuation) region is denoted by $\mathcal{S}_{t}$ (resp. $\mathcal{C}_{t}$). That is, for $t\in[0,\,T)$,
Although the shape of the surrender region can vary significantly (see for example Figures 4 and 5 of MacKay et al. (2017), when the fee rate is constant and the (time-dependent) surrender charge function satisfies certain conditions, MacKay (2014), Appendix 2.A shows that the surrender region is a connected set. The first part of Theorem 4.3 below indicates that as soon as the surrender region is non-empty and the fee and surrender charge functions depend only on time, then $t$-section of $\mathcal{S}$ has the form $\mathcal{S}_{t}=[b(t),\infty)$ for some $b(t)\in\mathbb{R}_{+}\cup\{\infty\}$. This result builds on Jacka (1991), Proposition 2.1 in the context of a bounded and time-homogeneous reward function in a time-homogeneous market model.
The second part of Theorem 4.3 confirms that the optimal stopping time obtained in Proposition 3.1 is unique.
Theorem 4.3.
Let $L$ be defined by (18).
  1. (i)
    If $L(t)<0$ for all $t\in[0,T)$ then the $t$-sections of $\mathcal{S}$, denoted by $\mathcal{S}_{t}$, are of the form
  2. (ii)
    If $L(t)\geq 0$ for all $(t)\in[0,T)$, then $\mathcal{S}=\emptyset$ and $T$ is the unique optimal stopping time for (7).
The proof of Theorem 4.3 is presented at the end of the section, as it builds upon additional results presented below.
@@ 4.4. Characterization of the Surrender Region
The proof of part (ii) of Theorem 4.3 does not require Assumptions 4.1 and 4.2 to hold. In this more general setting, if $L(t,x)\geq 0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, then $\mathcal{S}=\emptyset$ and $T$ is the unique optimal stopping time for (7).
Under the assumptions of Theorem 4.3, $b(t)$ is the smallest account value for which it is optimal to surrender the contract at time $t<T$, and for any account value greater than $b(t)$, it is also optimal to surrender, so that
and
The rest of this section presents further results on the (non-)emptiness of the surrender region and ends with the proof of Theorem 4.3. Many of the results presented here are inspired by the work of Villeneuve (1999), which we adapted to the time-dependent and discontinuous reward function considered in (7). As pointed out by Villeneuve (1999), Remark 2.1, adapting some of their results to a time-dependent payoff is not trivial.
Proposition 4.6 characterizes the non-emptiness of the surrender region. The first part involves $\text{relint}(\mathcal{S}_{t})$, the relative interior of a $t$-section $\mathcal{S}_{t}$, which is defined as its interior within the affine hull of $\mathcal{S}_{t}$. If it is not empty, the relative interior of $\mathcal{S}_{t}$ forms a vertical line in the plane. If $\text{relint}(\mathcal{S}_{t})\in\text{int}(\mathcal{S})$, it contains some finite $x$ for which $(t,x)\in\text{int}(\mathcal{S})$. The second part of Proposition 4.6 (ii) can be seen as a local version of the first part.
@@
    Let $\mathcal{S}\neq\emptyset$. For each $t\in[0,T)$, if $\text{relint}(\mathcal{S}_{t})\subset\text{int}(\mathcal{S})$, then $L(t)\leq 0.$
@@
    If $L(t)<0$ for all $t\in[0,T)$, then
    • $\bullet$
      $\mathcal{S}\neq\emptyset$ and $\mathcal{C}\neq\emptyset$, and
    • $\bullet$
      $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$.
@@
    From Theorem 4.2, Remark 4.9 and Corollary 4.2, we have that $v\in C^{0,1}([0,T]\times\mathbb{R}_{+}\setminus\{T\}\times\{G\})\cap C^{1,2}([0,T]\times\mathbb{R}_{+}\setminus\partial\mathcal{C})$, so that $v\in C^{1,2}(\text{int}(\mathcal{S}))$. Observe that for $x\in\text{relint}(\mathcal{S}_{t})$, $v(t,x)=xg(t)$, so that $xL(t)=\mathcal{L}_{t}v+v_{t}-rv$. We note that the differential operator can be applied since $x\in\int(\mathcal{S})$. Recall from Proposition 4.2 that under Assumption 4.2, $\mathcal{L}_{t}v+v_{t}-rv\leq 0$ for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$. The result follows since $x>0$ in $\text{int}(\mathcal{S})$.
@@
    To show the first statement, we use the continuation premium representation (28), so that
@@
    for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. We want to show that there exists some $(t_{1},x_{1})\in[0,T)\times\mathbb{R}_{+}$ satisfying $v(t_{1},x_{1})=\varphi(t_{1},x_{1})$ and some $(t_{2},x_{2})\in[0,T)\times\mathbb{R}_{+}$ satisfying ${v(t_{2},x_{2})>\varphi(t_{2},x_{2})}$. We proceed by contradiction.
    First, suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{S}=\emptyset$. Fix $t\in[0,T)$. Since $\mathcal{S}=\emptyset$, we have $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, so that (41) becomes
@@
    Observe that $L(s)=g_{t}(s)-c(s)g(s)<0\Leftrightarrow\frac{\,\mathrm{d}\ln g(s)}{\,\mathrm{d}s}<c(s)$ for all $s\in[0,T)$. By integrating from $t$ to $T$ on both sides of the inequality and using the fact that $g(T)=1$, we find that
@@
    To show that $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $[t_{1},t_{2})\subset[0,T)$, we proceed by contradiction. Fix $0\leq t_{1}<t_{2}<T$. Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\cup_{t_{1}\leq t<t_{2}}\mathcal{S}_{t}=\emptyset$ This implies $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, so that $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Using the continuation premium representation (28) with $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, we have
    $\begin{split}v(t,x)-\varphi(t,x)&=\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\\ &\quad+\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\right]\,\mathrm{d}s>0,\end{split}$
    so that,
    $\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]$ $>-\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\right]\,\mathrm{d}s$
    (45) $=-x\sup_{t\leq r<t_{2}}L(r)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s$
    Note that the function on the right-hand side is positive, equal to 0 when $x=0$, strictly increasing in $x$ (since $L(t)<0$ by assumption) and approaches infinity as $x\rightarrow\infty$. Then, using arguments similar to those of the proof of Proposition 4.6 (ii), there must exist ${y\in\mathbb{R}_{+}}$ such that for all ${x\geq y}$, ${\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\leq-x\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s}$, contradicting the strict inequality in (45) which must hold for all ${x\in\mathbb{R}_{+}}$. We conclude that or $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$.
Proposition 4.7 characterizes the emptiness of the surrender region.
Proposition 4.7.
Let $L$ be defined by (18).
  1. (i)
    For each $t\in[0,T)$, if $L(t)>0$, then $\mathcal{S}_{t}=\emptyset$ or $\mathcal{S}_{t}\in\partial\mathcal{S}$.
  2. (ii)
    For any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t)>0$ for all $(t)\in[t_{1},t_{2})$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
  3. (iii)
    If $\mathcal{S}=\emptyset$, then $g(t)-e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}\leq 0$ for all $t\in[0,T)$.
Remark 4.16.
Some of the assumptions of Proposition 4.7 can be relaxed. The proof of part (i) only requires Assumption 4.2. Furthermore, Proposition 4.7(ii) holds in a general setting, without Assumptions 4.1 and 4.2. In this general setting, it holds that for any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
@@
    This statement is obtained by observing the contrapositive of Proposition 4.6 (i).
@@
    Observe that $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$ implies ${\mathcal{C}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=[t_{1},t_{2})\times\mathbb{R}_{+}}$. Thus, we need to show that $v(t,x)>\varphi(t,x)$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$ and recall from the proof of Lemma 4.3 that for any $0<t<s\leq T$
@@
    for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, which completes the proof.
  3. (iii)
    Since $\mathcal{S}=\emptyset$, $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Then, for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ implies
    (47) $g(t)-e^{-\int_{t}^{T}c(u)\,\mathrm{d}u}\leq 0$
    for all $t\in[0,T)$.
Remark 4.17.
The first statements of Propositions 4.6 and 4.7 can apply to surrender charge functions that depend on both time and the value of the underlying process. Indeed, the proof relies on the results of Proposition 4.2 which only requires the fee function, and not the surrender charge, to depend exclusively on time, see Remark 4.10 for details.